OpenStudy (anonymous):

A group of 3 children an 2 adults pay a total of $120 to take a karate class. A group of 5 children and 1 adult take the same karate class for $95. What is the total cost for 1 child and 1 adult to take the karate class?

OpenStudy (goformit100):

Take unitary method in use to solve this Question.

OpenStudy (anonymous):

what? i dont know wh that is.

OpenStudy (anonymous):

this isan algebra 1 problem

OpenStudy (goformit100):

@electrokid Plz Explain This Sir

OpenStudy (anonymous):

\[3c+2a=120\\5c+a=95\] solve the equations

OpenStudy (anonymous):

c = price for 1 child a = price for 1 adult

OpenStudy (anonymous):

the question asks for \[a+c=?\]